what is the order of

i) 14 in Z24/ _

<8>?

ii) (Z10⊕U(10)/<2,9>?


1
Expert's answer
2021-06-15T13:54:32-0400

i)

(8) has order 3 in Z24 , so Z24/(8) is cyclic of order 8, generated by [14] +(8).

Now (8) = {[0], [8], [16]} ⊂ Z24

[14] +(8)"\ne" (8) (since 14 "\notin" (8))

([14] + (8))2 = [28] +(8)=[4] +(8) "\ne" (8)

([14] + (8))3 = [42] +(8)=[18] +(8) "\ne" (8)

([14] + (8))4 = [56] +(8)=[8] +(8) "=" (8)

([14] + (8))5 = [70] +(8)=[22] +(8) "\ne" (8)

([14] + (8))6 = [84] +(8)=[12] +(8) "\ne" (8)

([14] + (8))7 = [98] +(8)=[2] +(8) "\ne" (8)

([14] + (8))8 = [112] +(8)=[16] +(8) "=" (8)

Therefore, [14] + (8) has order 6 in Z24/(8)


ii)

"\Z^{\times}_{10}=\{1,3,7,9\}"

"\Z_{10}=\{0,1,2,3,4,5,6,7,8,9\}"

"<2,9> = (2,9),(4,1),(6,9),(8,1),(0,9),(2,1),(4,9),(6,1),(8,0),(0,1)"


The order of quotient group "\Z_{10}\times\Z_{10}^{\times}/<2,9>" :

"\frac{|\Z_{10}\times\Z_{10}^{\times}|}{|<2,9>|}=\frac{10}{10}=1"